sue can sholver snow from her driveway in 40 minutes. tom can do the same job in 50 minutes. How long would it take sue and tom to shovel the driveway if they worked together?
Answer by stanbon(75887)

sue can shovel snow from her driveway in 40 minutes. tom can do the same job in 50 minutes. How long would it take sue and tom to shovel the driveway if they worked together.
---
Sue rate = 1/40 job/min
Tom rate = 1/50 job/min
---------------
Together rate = 1/x job/min
---------------------------------
Equation:
rate + rate = together rate
1/40 + 1/50 = 1/x
50x + 40x = 40*50
90x = 2000
x = 22.222 minutes (time to do the job together)
--------------------------
Cheers,
Stan H.